A machine is set to pump cleanser into a process at the rate of 10 gallons per minute. Upon inspection, it is
learned that the machine actually pumps cleanser at a rate described by the uniform distribution over the
interval 9.5 to 12.5 gallons per minute. Find the probability that between 10.0 gallons and 11.0 gallons are
pumped during a randomly selected minute.
The answer is 0.33, how do I arrive at that answer?

Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

A machine is set to pump cleanser into a process at the rate of 10 gallons per minute. Upon inspection, it is
learned that the machine actually pumps cleanser at a rate described by the uniform distribution over the interval 9.5 to 12.5 gallons per minute.
------
Mean = 10---
6*sigma = 12.5-9.5 = 3
---
sigma = 1/2
-----
Find the probability that between 10.0 gallons and 11.0 gallons are
pumped during a randomly selected minute.
---
z(11) = (11-10)/3 = 1/3
z(10) = (10-10)/3 = 0
---
P(10< x < 11) = P(0< z < 1/3) = 0.1306
=====
I don't agree with the answer you were given.
=========================
Cheers,
Stan H.
=============
The answer is 0.33, how do I arrive at that answer?

Answer by edjones(8007)

s=1/2=.5
z=(11-10)/.5=2
Probability = .4772
.
Ed